OpenStudy (anonymous):

Could someone explain this? e^(2*S1*L)= -1 S1=((2n+1)*pi*i)/(2L)

OpenStudy (anonymous):

\[e ^{2 * s1*L}=-1\] \[s1 = ((2n+1)\pi*i)/(2L)\]

OpenStudy (ash2326):

put s1 in the given equation e^(2*((2n+1)pi*i)/2L*L) let's just simplify the term in power we have 2*((2n+1)pi*i)*L/2L 2L in numerator will get cancle with 2L in denominator , we 're left with (2n+1)pi*i if we put n=0 we het pi*i , if we put n=1 we get 3pi*i and so on we e^((2n+1)i) we know \[e^(\theta*i)=\cos \theta+i \sin \theta\] we know cos pi=-1 and sin pi=0 and so on cos 3pi=-1 sin 3pi=0 therfore our equation reduces to cos (2n+1)pi+isin (2n+1)pi cos(2n+1)pi=-1 sin(2n+1)pi=0 therfore we get -1

OpenStudy (anonymous):

uehhh, but what is you didnt know the answer... of S1 how would you solve it for S1?

OpenStudy (ash2326):

i'm not using the answer, i'm putting s1's value in equation and then reducing it. it's a standard equation cos(2n+1)pi=-1 sin(2n+1)pi=0

OpenStudy (ash2326):

oh you mean we have to find s1 sorry that's also easy

OpenStudy (anonymous):

yeahh i mean... how can i find s1? and get the answer what i have typed above :)

OpenStudy (ash2326):

we are given that e^(2*S1*L)=-1 that means that term in power is an odd multiple of pi so 2*S1*L=(2n+1)pi so S1=(2n+1)pi/2L

OpenStudy (ash2326):

where n=0,1,2,3

OpenStudy (anonymous):

yeah but why is it odd?

OpenStudy (anonymous):

how do you see that this equation is odd?

OpenStudy (ash2326):

because odd multiple of pi gives cos pi=-1 and sin pi=0

OpenStudy (ash2326):

as there is no i term in the result , that means sin term is 0 which can only happen when theta = 0,pi or 2pi but we want -1 cos pi , cos 3pi , cos5pi=-1 only odd multiple result in -1 therfore in general the theta is (2n+1)pi where n=0,1,2 if u put n=0 we get pi , n=1 3pi and so on

OpenStudy (zarkon):

n can also be negative

OpenStudy (anonymous):

you typed: we are given that e^(2*S1*L)=-1 that means that term in power is an odd multiple of pi so 2*S1*L=(2n+1)pi so S1=(2n+1)pi/2L The answer is S1=(2n+1)pi/2L * i

OpenStudy (anonymous):

where did your "i" go

OpenStudy (ash2326):

i will come

OpenStudy (ash2326):

only the term other than i is equal to odd multiple of pi

OpenStudy (ash2326):

i missed it but e^i*theta =cos theta +i sin theta

OpenStudy (anonymous):

i still dont get is :(

OpenStudy (anonymous):

could someone draw it out? i'm sure it has to do something with the unit circle

OpenStudy (jamesj):

\[ e^z = -1 \] if and only if \[ z = \pi i + 2\pi k i, \ \ k \in \mathbb{Z} \] Do you understand this much?

OpenStudy (anonymous):

hi james, no not really :s, but it consist only the Im. component

OpenStudy (jamesj):

If we write z = a + bi, then \[ e^z \] equals what?

OpenStudy (jamesj):

where a and b are real

OpenStudy (anonymous):

a is the Re b is the Im

OpenStudy (jamesj):

of z. Yes. Now what is e^z equal to?

OpenStudy (anonymous):

its must lie in the unit circle so i guess +1 or -1

OpenStudy (jamesj):

ok. What does \[ e^{i \theta } \] equal?

OpenStudy (anonymous):

i really dont know james...

OpenStudy (jamesj):

For what value of theta does e^(i . theta) = 1 ?

OpenStudy (anonymous):

owhh... if theta is 0 or pi/2 = 1 if theta is pi or 3/2pi = -1 am i right?

OpenStudy (anonymous):